# Memory Gym: Partially Observable Challenges to Memory-Based Agents in Endless Episodes

### v1

@@ 4.2 Endless Environments: GRU is more effective than Transformer-XL
Across all three endless environments, the recurrent agent (GRU) consistently outperforms the transformer-based agent (TrXL) by a significant margin. Notably, the most substantial performance gap between GRU and TrXL emerges in the results of Endless Mortar Mayhem (Figure 11(a)). Within this context, GRU attains an impressive IQM of 120 executed commands using roughly 300 million steps, whereas TrXL only attains an IQM of 18 commands. Furthermore, the incorporation of observation reconstruction exacerbates this performance difference, leading TrXL to a further deterioration, resulting in an IQM of 10. Furthermore, GRU exhibits consistently longer IQM episode lengths, with TrXL achieving an IQM of 288 steps and GRU reaching a peak IQM episode length of 3462 steps.
The outcomes depicted in Figure 11(b) illustrate the results obtained from the Endless Mystery Path environment. In this case, GRU remains more effective than TrXL, although the gap between their performances is narrower compared to EMM. GRU achieves an IQM return of 2.65, with TrXL achieving 2.02. Although neither memory approach converges within the allocated training budget of 819 million steps, GRU displays a more pronounced upward trend in performance. In terms of episode length, neither agent achieves lengths as seen in EMM. GRU attains an IQM episode length of 600 steps in EMP, while TrXL reaches up to 400 steps. This divergence can be attributed to the episode’s dynamics, where agents tend to fall off after significant progress, necessitating more steps to regain lost ground and accumulate further rewards.
@@ 4.3 Investigating Transformer-XL’s Surprisingly Low Effectiveness
Figure 12: Experiments on Endless Mortar Mayhem with varied TrXL configurations. The learning rate schedule (LR) is adjusted to decay from $42.75e-$ 4 to $41.0e-$ 4 over 160 million steps, compared to the previous $51.0e-$ 5. We also incorporate observation reconstruction (Obs. Rec.) and observe the most significant improvement upon introducing a ground truth estimation head (GT) as sanity check. Another test augments TrXL’s query with relative positional encoding (QPos). The final, but inferior, experiment combines the optimized learning rate, augmented query, and ground truth estimation.
The unexpectedly low performance of TrXL prompts us to several hypothesis that are explored within the next subsections. Endless Mortar Mayhem poses the largest performance gap and therefore we utilize this environment for further investigations.
@@ 4.3.1 Inadequate Networks Capacity
Our TrXL baseline comprises 2.8 million trainable parameters, while GRU consists of 4.05 million, prompting the question of whether TrXL’s model architecture lacks the necessary capacity. To investigate this, we conducted experiments varying the number of blocks (2, 3, 4), the embedding dimension size (256, 384, 512), and the memory window length (256, 384, 512). None of these adjustments led to closing the performance gap to GRU.
Scaling up TrXL also increases its demand for GPU memory. When scaling up multiple architecture details, our implementation and hyperparameters may cause the training to exceed the available 40GB GPU memory of an NVIDIA A100. Workarounds include reducing the batch size or transferring training data between the GPU and CPU, but these options significantly worsen wall-time efficiency.
Furthermore, another indication of untapped capacity lies in the amplification of the learning signal, which we elaborate on in the following section.
@@ 4.3.2 Weak Learning Signal
Figure 12 depicts several additional experiments aimed at amplifying the learning signal. We made naive a adjustment to the learning rate schedule, utilized observation reconstruction, and introduced a ground truth estimation head. This ground truth estimation head is added to the actor-critic model architecture and predicts the target position to move to. Labels are provided by the environment, and the estimates are optimized using the mean-squared error loss. While ground truth information is typically unavailable and considered inappropriate for this benchmark, it serves as a useful sanity check. It helps determine if an additional learning signal is advantageous in this context. When training incorporates ground truth estimation, there is a noticeable improvement in the agent’s policy. Previously, the IQM of completed commands stood at 18; with ground truth estimation, it reaches 50. This outcome implies that the model’s capacity might not be the primary constraint. However, using the ground truth estimation head can lead to instability in training, with the agent’s performance fluctuating between 20 and 50 completed commands after about 300 million steps.
During the initial 160 million training steps, the learning rate linearly decays from $2.75$e-$4$ to $1.0$e-$5$. By setting the final decayed learning rate to $1.0$e-$4$, the IQM performance hits 24 commands. Lower performances are observed when adding the observation reconstruction head to either just TrXL or TrXL with the optimized learning rate schedule.
In summary, augmenting the learning signal — either by adjusting the learning rate or by incorporating ground truth estimation — leads to more performant policies.
@@ 4.3.3 Lack of Temporal Information in the Initial Query
The next hypothesis pertains to the initial query of the first TrXL block. This query is constructed solely based on the features extracted from the observation encoders at the current time step, encompassing only minimal temporal information. In contrast, the query of the second block draws from the aggregated outcome of the memory window, thus capturing more substantial temporal information. However, we believe that the initial query could be further enriched with additional information. In Figure 12, one experiment augments the query with relative positional encoding, providing direct access to the current time step. Despite this enhancement, the agent’s performance only moderately improves, reaching an IQM of 19 completed commands. We also explored embedding the original query with the previous output of the last TrXL block. However, this simple approach did not yield positive results.
All of the aforementioned measures by far do not reach the level of the GRU agent. Even if combined, a low performance of 33 executed commands is achieved, which is also relatively unstable.
@@ 4.3.4 Harmful Off-Policy Data in the Episodic Memory
At last it can be discussed whether the combination of PPO and TrXL’s episodic memory pose a threat. PPO is an on-policy algorithm that consumes the training data for a few epochs. Once the second training epoch on the same batch commences, the batch is already considered off-policy. However, PPO’s loss function can mitigate this issue and hence allows for several epochs. But this mitigation is likely limited and we wonder whether this issue is enlarged by the episodic memory. In our current training setup, an agent collects 512 steps of data. If an episode exceeds 512 steps and is still running, the earlier steps from that episode become outdated over several epochs. In future work, it needs to be investigated if a stale episodic memory, which is based on off-policy data, hinders training progress. Frequently refreshing the episodic memory is expected to be expensive in terms of wall-time.
5 Related Work
We begin by describing different approaches aimed at improving agents’ capabilities through memory mechanisms, and highlight the challenge of accessing related baselines. Subsequently, we present an overview of existing memory benchmarks, comparing them conceptually to Memory Gym.
5.1 Memory-based Approaches
The work of Mnih et al. (2015) on mastering Atari games (Bellemare et al., 2013) using Deep Q-Networks (DQN) marked a significant milestone in the advancement of deep reinforcement learning algorithms. To capture temporal information, such as velocities and accelerations, a frame stacking approach was employed. More advanced memory-based approaches were introduced, which we will briefly detail subsequently.
Deep Recurrent Q-Networks (DRQN) (Hausknecht and Stone, 2015) combine DQN with an LSTM. The replay buffer stores sequences, while initializing the recurrent hidden state to zero during training. The effectiveness of this approach is demonstrated on Atari games that involve frame flickering.
Recurrent Replay Distributed DQN (R2D2) (Kapturowski et al., 2019) extends DRQN with a distributed prioritized experience replay buffer, storing overlapping sequences of fixed length. A burn-in period is utilized to address staleness in recurrent hidden states by propagating a portion of the sequence for a more effective starting state. R2D2’s performance is evaluated on Atari games (Bellemare et al., 2013) and DMLab-30 (Beattie et al., 2016).
The Memory Recall Agent (MRA) (Fortunato et al., 2019), built on IMPALA (Espeholt et al., 2018), employs a two-fold memory mechanism: a working memory based on an LSTM and a slot-based episodic memory containing summaries of past experiences. By utilizing a query-key approach, information is read from and written to the episodic memory. MRA is evaluated on the Memory Task Suite contributed by the same publication.
Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) adds the concept of Transformer-XL (TrXL) to enable an agent, trained with V-MPO (Song et al., 2020), to leverage memory. By incorporating an identity map reordering and a gating mechanism inspired by GRU, TrXL is further improved. GTrXL’s performance is demonstrated on DMLab-30 (Beattie et al., 2016), Numpad, and Memory Maze.
The Hierarchical Chunk Attention Mechanism (HCAM) (Lampinen et al., 2021) builds upon TrXL by partitioning the agent’s memory into chunks and selecting the most relevant chunks based on their summaries to attend to their stored information in detail. HCAM is trained using IMPALA (Espeholt et al., 2018) and evaluated on multiple environments including Dancing the Ballet, Object Permanence, Rapid Word Learning, Visual Match, Paired Associative Inference, and One-Shot StreetLearn Navigation.
History comprEssion via Language Models (HELM) (Paischer et al., 2022a) utilizes pre-trained language models to compress a history of observations into a single context vector (memory). It employs a frozen TrXL encoder, pre-trained on the WikiText 103 dataset (Merity et al., 2017). HELM is evaluated on the Procgen (Cobbe et al., 2020) and Minigrid (Chevalier-Boisvert et al., 2018) environments.
5.2 Lack of Accessible Memory Baselines
The accessibility of the approaches to memory described above poses a common issue. While HELM is fully open source and successfully integrated into our training framework, the implementations of DRQN, R2D2, MRA, and GTrXL have not been made publicly available by their original authors. HCAM is partially open source regarding its model architecture. The lack of open source implementations makes reproducing their results more challenging, as implementation details play a crucial role. This issue is particularly pronounced for GTrXL, as to the best of our knowledge, no one has yet been able to reproduce its underlying DRL algorithm, V-MPO.
Within the community, RLlib (Liang et al., 2018), DI-engine (DI-engine Contributors, 2021), and Brain Agent (Lee et al., 2022) advertise to feature GTrXL. However, these frameworks are large and complex, supporting multiple DRL algorithms, model architectures, and environments, which hampers verification, debugging, gaining insights, and making modifications. We tested RLlib with Memory Gym, but obtained negative returns even though the environments do not signal negative rewards. DI-Engine’s GTrXL implementation is based on R2D2, and our experiments revealed poor sample throughput by a magnitude of 10 compared to our GRU baseline. When debugging Brain Agent’s implementation, we noticed an unexpected discrepancy in the query length between data sampling and optimization. While we do not claim that these frameworks are dysfunctional, we present our limited experience in exploring them. We believe that our contributed open source baselines are easy to follow, which facilitates reproducibility and further research in the field.
5.3 Related Reinforcement Learning Memory Benchmarks
Previous studies have explored the use of memory-based agents in a variety of environments. Some of them are originally fully observable but are turned into partially observable Markov Decision Processes (POMDP) by adding noise or masking out information from the agent’s observation space. For instance, this was done for the Arcade Learning Environment (Bellemare et al., 2013) by using flickering frames (Hausknecht and Stone, 2015) and common control tasks by removing the velocity from the agent’s observation (Heess et al., 2015; Meng et al., 2021; Shang et al., 2021).
Control tasks also touch on the context of Meta-Reinforcement Learning, where memory mechanisms are prominent (Wang et al., 2021; Melo, 2022; Ni et al., 2022). The same applies to Multi-Agent Reinforcement Learning (Berner et al., 2019; Baker et al., 2020; Vinyals et al., 2019). As we solely focus on benchmarking the agent’s memory and its ability to generalize, we do not compare Memory Gym to environments of more complex contexts such as DM Alchemy (Wang et al., 2021), Crafter (Hafner, 2021), or Obstacle Tower (Juliani et al., 2019). These might need additional components to the agent’s architecture and its training paradigm, for instance to approach a notable exploration challenge.
When skimming the approaches described in section 5.1, it appears that there is a lack of consensus regarding the choice of environments for evaluation as they assess their approaches on different benchmarks, thereby hindering comparisons between them. Reasons for this observation are either missing accessibility, slow simulation speeds, or weak memory challenges (Pleines et al., 2023). Subsequently we give a coarse overview of recently used benchmarks chronologically.
Deepmind Lab 30 (Beattie et al., 2016) presents a collection of 30 procedurally generated first-person 3D environments. Starting from a general motor-control navigation task and visual observations, each environment poses a different goal such as collecting fruit, playing laser tag or traversing a maze. Parisotto et al. (2020) categorize these environments into memory and reactive tasks.
PopGym (Morad et al., 2023) has been published at ICLR 2023 like our original work contributing Memory Gym (Pleines et al., 2023). It is a benchmark consisting of 15 environments that are tagged as diagnostic, control, noisy, game, or navigation. These environments are designed to quickly converge by providing vector observations instead of visual ones.
5.4 Comparing Memory Gym to Related Benchmarks
Memory Gym’s environments offer a unique feature that sets them apart from related environments: an endless design tailored to rigorously test memory capabilities. It is crucial to differentiate between “endless” in this context and “open-ended” environments, such as Minecraft (Fan et al., 2022). While open-ended environments aim for agents to autonomously acquire a broad range of diverse and general skills without human oversight, Memory Gym’s ”endless” design remains anchored to a specific task structure.
The finite environments Mortar Mayhem and Mystery Path exhibit both similarities and unique characteristics compared to related environments. Several environments, such as Ballet, PopGym’s Autoencode, Spot the Difference, and Minigrid Memory, require the agent to initially memorize at least one cue to solve the task. While Ballet, Spot the Difference, and Minigrid Memory involve mapping clues to a single solution, Mortar Mayhem and PopGym’s Autoencode challenge the agent to solve a sequential puzzle. In Mortar Mayhem, delays between showing cues and executing commands can be dynamically sampled, similar to the concept in Ballet. Mystery Path is most similar to the navigation concepts of the environments that are inspired by the Morris water maze. Once fallen off, the agent must return to the point before where it lastly fell off and then has to take different actions to make progress on the hidden path. Concerning Searing Spotlights we were unable to identify an already existing environment that shares conceptual details.

### v3

@@ 5.2 Endless Environments: GRU is more effective than Transformer-XL
To our surprise, across all three endless environments, the recurrent agent consistently outperforms the transformer-based agent by a significant margin. This diverges from the observations made by Parisotto et al. (2020) and Lampinen et al. (2021), where an LSTM was found to be less effective. Notably, the most substantial performance gap between GRU and TrXL emerges in the results of Endless Mortar Mayhem (Figure 7(a)). Within this context, GRU attains an impressive IQM of 125 executed commands using roughly 300 million steps, whereas TrXL only attains an IQM of 18 commands. Furthermore, the incorporation of observation reconstruction exacerbates this performance difference, leading TrXL to a further deterioration, resulting in an IQM of 10. Furthermore, GRU exhibits consistently longer IQM episode lengths, with TrXL achieving an IQM of 288 steps and GRU reaching a peak IQM episode length of 3462 steps.
The outcomes depicted in Figure 7(b) illustrate the results obtained from the Endless Mystery Path environment. In this case, GRU remains more effective than TrXL, although the gap between their performances is narrower compared to Endless Mortar Mayhem. GRU achieves an IQM undiscounted return of 2.65, with TrXL achieving 2.02. Although neither memory approach converges within the allocated training budget of 819 million steps, GRU displays a more pronounced upward trend in performance. In terms of episode length, neither agent achieves lengths as seen in in Endless Mortar Mayhem. GRU attains an IQM episode length of 600 steps in Endless Mystery Path, while TrXL reaches up to 400 steps. This divergence can be attributed to the episode’s dynamics, where agents tend to fall off after significant progress, necessitating more steps to regain lost ground and accumulate further rewards.
@@ 5.3 Investigating Transformer-XL’s Surprisingly Low Effectiveness
Figure 8: Experiments on Endless Mortar Mayhem with varied TrXL configurations. The learning rate (LR) is adjusted to decay from $42.75e-$ 4 to $41.0e-$ 4 over 160 million steps, compared to the previous $51.0e-$ 5. We also incorporate observation reconstruction (Obs. Rec.) and observe the greatest improvement based on a ground truth estimation head (GT) as sanity check. Another test augments TrXL’s query with relative positional encoding (QPos). The final experiment combines the optimized learning rate, augmented query, and ground truth estimation.
The unexpectedly low performance of TrXL prompts us to examine several hypotheses. Endless Mortar Mayhem, presenting the largest performance gap, serves as our primary environment for these detailed investigations.
@@ 5.3.1 Inadequate Networks Capacity
Our TrXL baseline comprises 2.8 million trainable parameters, while GRU consists of 4.05 million, prompting the question of whether TrXL’s model architecture lacks the necessary capacity. To investigate this, we conducted experiments varying the number of blocks (2, 3, 4), the embedding dimension size (256, 384, 512), and the memory window length (256, 384, 512). None of these adjustments led to closing the performance gap to GRU. Scaling up TrXL also increases its demand for GPU memory. When scaling up multiple architecture details, our implementation and hyperparameters may cause the training to exceed the available 40GB GPU memory of an NVIDIA A100. Workarounds include reducing the batch size or transferring training data between the GPU and CPU, but these options significantly worsen wall-time efficiency. Furthermore, another indication of untapped capacity lies in the amplification of the learning signal, which we elaborate on in the following section.
@@ 5.3.2 Weak Learning Signal
Figure 8 depicts several additional experiments aimed at amplifying the learning signal. We made a naive adjustment to the learning rate schedule, utilized observation reconstruction, and introduced a ground truth estimation head. This ground truth estimation head is added to the actor-critic model architecture and predicts the target position to move to. Labels are provided by the environment, and the estimates are optimized using the mean-squared error loss. While ground truth information is typically unavailable and considered inappropriate for this benchmark, it serves as a useful sanity check. It helps determine if an additional learning signal is advantageous in this context. When training incorporates ground truth estimation, there is a noticeable improvement in the agent’s policy. Previously, the IQM of completed commands stood at 18; with ground truth estimation, it reaches 50. This outcome implies that the model’s capacity might not be the primary constraint. However, using the ground truth estimation head can lead to instability in training, with the agent’s performance fluctuating between 20 and 50 completed commands after about 300 million steps. During the initial 160 million training steps, the learning rate linearly decays from $2.75$e-$4$ to $1.0$e-$5$. By setting the final decayed learning rate to $1.0$e-$4$, the IQM performance hits 24 commands. Lower performances are observed when adding the observation reconstruction head to either just TrXL or TrXL with the optimized learning rate schedule. In summary, augmenting the learning signal — either by adjusting the learning rate or by incorporating ground truth estimation — leads to more performant policies.
@@ 5.3.3 Lack of Temporal Information in the Initial Query
The next hypothesis pertains to the initial query of the first TrXL block. This query is constructed solely based on the features extracted from the observation encoders at the current time step, encompassing only minimal temporal information. In contrast, the query of the second block draws from the aggregated outcome of the memory window, thus capturing more substantial temporal information. However, we believe that the initial query could be further enriched with additional information. In Figure 8, one experiment augments the query with absolute positional encoding, providing direct access to the current time step. Despite this enhancement, the agent’s performance only moderately improves, reaching an IQM of 19 completed commands. We also explored embedding the original query with the previous output of the last TrXL block. However, this simple approach did not yield positive results.
@@ 5.3.4 Harmful Off-Policy Data in the Episodic Memory
Next, it can be discussed whether the combination of PPO and TrXL’s episodic memory pose a threat. PPO is an on-policy algorithm that consumes the training data for a few epochs. Once the second training epoch on the same batch commences, the batch is already considered off-policy. However, PPO’s loss function can mitigate this issue and hence allows for several epochs. But this mitigation is likely limited and we wonder whether this issue is enlarged by the episodic memory. In our current training setup, the agent collects 512 steps of data. If an episode exceeds 512 steps and is still running, the earlier steps from that episode become outdated over multiple epochs. In future work, it needs to be investigated if a stale episodic memory, which is based on off-policy data, hinders training progress. Frequently refreshing hidden states, a process used in the off-policy algorithm R2D2 that relies on recurrence (Kapturowski et al., 2019), is expected to be computationally inefficient.
5.3.5 Indistinguishable Absolute Positional Encoding
The final hypothesis concerns a potential limitation of the absolute positional encoding. Without positional encoding, the agent cannot differentiate the time between past observations. As a default, we rely on absolute positional encodings. This encoding always spans across the potential maximum episode length. In the finite environments, the longest range is set to 512, being the sequence length, which is also used in the original transformer (Vaswani et al., 2017). For the endless tasks, this encoding is set up for 2048 steps, an episode duration that is yet unreachable to the TrXL agents.
The original work of TrXL introduces a relative positional encoding (Dai et al., 2019) as a consequence of extended context of lengths. However, when they refer to the absolute positional encoding, they only apply it to the range of the current window. If a window has a length of 4, the positions within the window are always the same (i.e. $[0,1,2,3]$), even though the window moves across the entire sequence. In our case, the absolute positional encoding considers the entire sequence. If the window has progressed in time, the applied positional encoding relies on the current time step (i.e. $[t-3,t-2,t-1,t]$).
Nevertheless, the absolute positional encoding needs to be defined prior training. When scaling it up, the question arises whether the agent’s model is capable of differentiating time steps, because a longer length indicates a more fine-grained sample frequency over the underlying positional encoding. To test this, we trained a TrXL agent on the finite task of Mortar Mayhem and configured an absolute positional encoding of length 1024 and 2048. Both experiments, which were repeated twice, resulted in an agent that poorly completes only 1 to 2 commands. This demands for questioning the absolute positional encoding in endless tasks. But if the elements of the absolute position encoding of length 2048 are not distinguishable, how do those agents master more than 20 commands in Endless Mortar Mayhem?
Figure 9: Experiments on Endless Mortar Mayhem with varied positional encoding. The learning rate schedule (LR) is adjusted to decay from $2.75$e-$4$ to $1.0$e-$4$ over 160 million steps, compared to the previous $1.0$e-$5$. The acronym LPE refers to a learned positional encoding, while RPE depicts the relative one.
If this assumption holds right that the absolute positional encoding is the culprit in our TrXL baseline, leveraging learned or relative positional encodings instead should improve performance. As seen in Figure 9, no notable advancement is accomplished by both variants. We observed two outliers among the 5 runs that utilized relative positional encoding. One underperformed, completing only 13 commands, while the other excelled, finishing more than 60 commands. Since the interquartile mean represents only the central 50% of the data, these outliers are not apparent in the sample efficiency curve. Given this discussion, a more in-depth exploration of positional encoding in this context is advisable.
5.4 Recurrence is not Vulnerable to Spotlight Perturbations
Training recurrent agents in Searing Spotlights with previously established hyperparameters (refer to the last column of Table 4) does not yield a meaningful policy. Despite numerous attempts at hyperparameter tuning, architecture adjustments, and task modifications, Searing Spotlights remained as a puzzle. Even with full observability, where all entities are permanently visible, the performance is catastrophic. These challenges prompted the hypothesis that spotlights might introduce noise, heavily affecting the agent’s policy (Pleines et al., 2023). Support for this theory came from experiments in a simplified BossFight environment (Cobbe et al., 2020), using identical spotlight dynamics as in Searing Spotlights, suggesting a potential vulnerability of recurrent agents to such perturbation (Pleines et al., 2023).
Table 1: Results from the ablation study on Searing Spotlights with full observability. Each experiment was run 5 times, though the data sourced directly from the training process deviates from our standard protocol. Every PPO iteration (of 5,000 total) produces a data point that reflects the average over the last 100 episodes. Rows are ordered in descending fashion by the mean success rate. The mean is computed from the top-performing data point, averaged across repetitions, from the specific experiment. Consequently, the time point is represented as a percentage duration. The gradient norm is accumulated over all data points within the given duration.
Experiment Value Success Rate Duration Gradient Norm
New Old Mean Std Mean Std
Advantage Norm. Off On 1.00 0.00 31% 0.02 0.01
No Memory 0.96 0.11 99% 0.28 0.03
Agent Health Points 5 10 0.79 0.35 99% 0.30 0.02
Transformer-XL 0.69 0.27 99% 0.30 0.03
Agent Speed 3.0 2.5 0.58 0.22 65% 0.33 0.06
Clip Range 0.1 0.2 0.52 0.25 99% 0.27 0.05
Max Episode Length 256 512 0.48 0.27 99% 0.37 0.04
Old Hyperparameters 0.43 0.06 96% 0.38 0.06
Observation Rec. On Off 0.39 0.06 96% 0.40 0.04
Max Gradient Norm 0.25 0.5 0.36 0.05 99% 0.24 0.01
However, the results presented in Figure 6(g) neglect this hypothesis. This prompted us to conduct an ablation study, beginning with the previous hyperparameters. In this study, only a single hyperparameter or environment option was replaced with the current one at a time. Table 1 presents the results for Searing Spotlights when full observability is ensured. Initially, observation reconstruction was presumed to be the solution. However, the crucial adjustment proved to be the decision to avoid normalizing advantage estimates during optimization. Only the GRU agent, which omits advantage normalization, and the memory-less agent, based entirely on previous hyperparameters, recorded noteworthy success rates of 100% and 96%, respectively.
A salient observation emerges when tracking the gradient norms across the agent models during training. With normalized advantages, the gradient norm is typically more than 12 times greater than when normalization is bypassed. This suggests that during gradient descent, excessively large optimization steps might be taken, resulting in an oscillating behavior that fails to converge to a local minimum. Simultaneously, unnormalized advantages remain at a scale of $2$e-$4$, but post-normalization, the magnitudes stretch beyond $5$. Considering that the policy ratio is modulated by these advantage estimates (Schulman et al., 2017), normalization introduces a strikingly different scale than its absence. Standard practice typically normalizes advantages by subtracting the mean and dividing by the standard deviation, often applied at the mini-batch level.
Several comprehensive studies have delved into such nuances of PPO’s implementation (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022b). Among them, only Andrychowicz et al. (2021) turned off advantage normalization and found negligible performance impact. However, their extensive study focused on control tasks with undiscounted returns in the thousands, contrasting Memory Gym’s finite environments where the undiscounted returns do not exceed 2. During hyperparameter optimization, turning off advantage normalization also enhanced performance in Mortar Mayhem and Mystery Path. Given the critical importance of utilizing raw advantages, there remains an opportunity to delve even deeper into PPO to comprehend and refine its essential aspects.